OpenStudy (anonymous):

(4-sq rt 3)(4+ sq rt 3) a. 13-8 sq rt 3 b. 5-8 sq rt 3 c. 13 d. 5

OpenStudy (anonymous):

hint use (a-b)(a+b)=a^2-b^2

OpenStudy (anonymous):

\[(4-\sqrt{3})(4+\sqrt{3})\]

OpenStudy (anonymous):

A ?

OpenStudy (anonymous):

no.... another hint \[(\sqrt{a})^{2}=a\]

OpenStudy (anonymous):

ok so b ?

OpenStudy (anonymous):

\[(4-\sqrt{3})(4-\sqrt{3})=4^{2}-(\sqrt{3})^{2}\] =\[16-3\] =\[13\]

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

the answer is C